OpenStudy (anonymous):

Find the center and the radius of the circle: x2 + 4x + y2 - 6x = 10.

OpenStudy (anonymous):

circle formula x^2+y^2=r^2 you need to put it in this form

OpenStudy (anonymous):

(x-h)^2+(y-k)^2=r^2 where h and k the shift from the (0,0))

OpenStudy (michele_laino):

@jorgev182 please if I simplify 4x-6x=-2x, I can write your formula like this: \[x^{2}+y ^{2}-2x=10\] is it right?

OpenStudy (anonymous):

yes thank you

OpenStudy (michele_laino):

if so, I can write your equation as@AJ01 said, namely as below: \[(x-1)^{2}+(y-0)^{2}=11\] from which you can read the coordinnate of the centre and the radius